# Robustness of contextuality under different types of noise as quantifiers for parity-oblivious multiplexing tasks

### v1

@@ Abstract
Generalised contextuality is a well-motivated notion of nonclassicality powering up a myriad of quantum tasks, among which is the celebrated case of a two-party information processing task where classical information is compressed in a quantum channel, the parity-oblivious multiplexing (POM) task. The success rate is the standard quantifier of resourcefulness for this task, while robustness-based quantifiers are as operationally motivated and have known general properties. In this work, we leverage analytical and numerical tools to estimate robustness of contextuality in POM scenarios under different types of noise. We conclude that for the 3-to-1 case robustness of contextuality to depolarisation, as well as a minimisation of robustness of contextuality to dephasing over all bases, are good quantifiers for the nonclassical advantage of this scenario. Moreover, we obtain a general relation between robustness of contextuality to depolarisation and the success rate in any $n$-to-1 POM scenario and show how it can be used to bound the number of bits encoded in this task.
@@
Since the advent of quantum technologies, the superiority of quantum over classical methods has been anticipated across various areas, such as computation, information processing, and metrology. A celebrated result by Holevo [1] suggests that communication tasks may be an exception, stating that one can never retrieve more than $\log{n}$ bits of classical information out of $n$ qubits, apparently forbidding quantum superiority over classical strategies for communicating. Quantum advantage, however, can arise in communication tasks that profit from the quantum systems’ capacity to carry more information than classical ones [2, 3]. Therefore, such tasks prove interesting from both practical and foundational perspectives: they point to advantageous technological opportunities using quantum systems and also showcase how Holevo’s theorem can be conciliated with such an advantage in communication. We here focus on one of such tasks, called parity-oblivious multiplexing (POM), in which Alice sends Bob an $n$-bit string by sharing a smaller number of qubits. It has been demonstrated that the maximum success probability using quantum strategies for POM tasks is strictly higher than the maximum success probability obtained using classical systems, constituting an instance of quantum advantage over classical strategies for communication [4].
@@
It was previously shown that quantum advantage in POM tasks is powered by contextuality [5], one of the best-motivated notions of nonclassicality available [6]. Besides POM, contextuality powers up other aspects of communication [7, 8, 9], as well as computation [10], machine learning [11], information processing [12, 13, 14, 15], metrology [16], state-dependent cloning [17], and state discrimination [18, 19, 20, 21]. Contextuality first appeared in the literature in the Kochen-Specker theorem [22], and only recently it has been preliminary employed to make statements about generalised probabilistic theories (GPTs) [23, 24, 25] other than quantum theory. Generalisations of this notion have been developed in recent years, and Spekkens introduced an alternative approach that extends the notion of contextuality to encompass preparations, transformations, and unsharp measurements [5], contrasting with the Kochen-Specker theorem that primarily focuses on projective measurements. Since then, generalised contextuality has been proved to subsume or be related to many common notions of nonclassicality [4, 26, 27, 10, 28, 18, 29, 30, 31, 32, 33], while challenging the nonclassical status of some phenomena [34, 35, 36, 37, 38]. Within the generalized contextuality framework – and therefore throughout this manuscript – a theory or scenario therein is deemed classical when it can be explained by a generalized noncontextual ontological model.
@@
Since we have good and operationally motivated measures for contextuality – robustness against different kinds of noise – and we know that contextuality powers up the advantage behind parity-oblivious multiplexing tasks, it is natural to ask how these contextuality measures relate to the natural quantifier in POM tasks, i.e., the success probability itself. In this work, we turn to this problem by exploring the generality allowed by the code in Ref. [39] to define different kinds of noise and compare its behaviour to the success probability in POM tasks. We find a general expression relating the robustness of contextuality to depolarisation and the success rate for $n$-to-1 POM tasks and show how it can be used to recover a well-known bound on the number of bits optimally encoded in a qubit in this task. Then, building up on Ref. [41], we generalise its modification to estimate robustness of contextuality to dephasing noise in a family of scenarios related to the 3-to-1 POM. Our numerical results show that robustness of contextuality to dephasing can also perform as a quantifier when minimised over specific axes. Moreover, since the code essentially tests for simplex-embeddability of noisy states and effects, the method we adopt here can also be used for testing a class of generalized probabilistic theories broader than just quantum theory.
@@ I.1 Noncontextuality
The so-called operational approach is a way to construct physical theories based on laboratory experience [23, 24, 43]. This means that instead of the physical properties of a system and their time evolution, operational theories focus on how this physical system can be prepared, manipulated, and investigated in general experimental setups. Therefore, this approach is broad in the sense that it can be used to describe different kinds of theories, quantum and classical theories included. A typical prepare-and-measure scenario is described in this framework by the possible preparation procedures $P\in\mathcal{P}$, measurement procedures $M\in\mathcal{M}$, and possible outcomes $k\in K$ for each measurement that can be implemented in the experiment, as well as the statistics $p(k|M,P)$ obtained by implementing these procedures. Moreover, these scenarios are also characterised by equivalences between different procedures: two preparations, denoted as $P$ and $P^{\prime}$, are deemed operationally equivalent, denoted as $P\simeq P^{\prime}$, if they cannot be distinguished empirically, even in principle. This implies that any possible measurement outcome yields the same probability for both preparations. Similarly, two measurement outcomes, denoted as $k|M$ and $k^{\prime}|M^{\prime}$, are considered operationally equivalent, denoted as $k|M\simeq k^{\prime}|M^{\prime}$, if they both yield the same probability for any possible preparation. The tuple $(\mathcal{P},\mathcal{M},K,p,\simeq)$ defines an operational theory. As previously mentioned, quantum theory is one possible operational theory.
@@
Generalized contextuality arises within this operational framework as the impossibility of an operational theory to conform with a noncontextual ontological model. An ontological model for an operational theory provides an underlying explanation for its statistics based on classical probability theory and Boolean logic, such that the behaviours captured by the operational theory can be reasoned about in classical terms. In particular, the ontological model maps the system to an ontic space $\Lambda$, preparation procedures are associated with probability distributions $\mu_{P}(\lambda)$, and measurement outcomes are associated with response functions $\xi_{k|M}(\lambda)$ over the ontic space, with $\lambda\in\Lambda$. This model explains the operational theory when, for all $P\in\mathcal{P}$ and $k|M\in K\times\mathcal{M}$, it satisfies
@@
Admitting of some ontological model is not yet a classical feature but a general one. Indeed, classicality has additional demands: any equivalences between operational procedures must hold at the ontological level [^2] Otherwise, further arguments must be provided on why the procedures are not operationally distinguishable, despite being ontologically different., which underpins the notion of noncontextuality [5, 6]. An ontological model satisfies the assumption of noncontextuality if, for any $P,P^{\prime}\in\mathcal{P}$, $k|M,k^{\prime}|M^{\prime}\in K\times\mathcal{M}$,
@@
The condition of admitting of some noncontextual ontological model has proven to be a precise definition of classicality since the noncontextual ontological model itself provides a classical explanation to the empirical predictions. As was shown before, quantum theory does not admit such explanations [5], featuring nonclassicality.
@@ I.2 Operational measures of contextuality
Often, assessing noncontextuality of an operational scenario requires the characterisation of a large ontic space [^3] In quantum scenarios, for instance, the ontic space might be as large as $|\Lambda|=(\text{dim}\mathcal{H})^{2}$ [39]. and the possible epistemic states and response functions over it that can explain the statistical data and conform to the assumption of nonlocality for all equivalence relations. To avoid these issues when assessing the nonclassicality of the examples investigated in this work, we will employ the linear program introduced in Ref. [39], which relies on results from the framework of generalised probabilistic theories (GPT) [23, 24, 25]. We will not enter into the details of how the operational scenario translates to the GPT description, but a summary of the linear program and the reasoning behind it is provided in Appendix A. The relevant aspect for the present work is that this linear program decides the existence of a simplex embedding for the GPT fragment (see Fig. 1), which in turn is equivalent to the existence of a noncontextual ontological model for the associated operational scenario [40]. If the program cannot find a simplex embedding for the input scenario, i.e., the scenario exhibits contextuality, it calculates how much depolarising noise $r$ is necessary so that the simplex embedding becomes possible. The implementation of the program takes in sets of states and effects of a GPT and outputs this minimal value of $r$ along with the respective noncontextual ontological model explaining the scenario partially depolarised by a factor $r$.
@@
The quantity $r$ hence may serve as an operational measure of nonclassicality, which we denote as the robustness of contextuality to depolarisation. It is known from simplex embedding reasoning that there is always a finite amount of partial depolarising noise under which an operational scenario becomes noncontextual, and explicit examples of this fact have been reported [18, 44, 45]. Alternatively, robustness can be defined in terms of different kinds of noise besides depolarising, so in this work, we also consider the case of dephasing. This is because it is a well-established fact that a completely dephased prepare-and-measure scenario will always admit of a noncontextual model [41], so the action of dephasing noise over a scenario will also eventually allow for a simplex embedding [^4] Although in this case, partial dephasing noise might not be enough to allow for the simplex embedding, as shown in Ref. [41].. Therefore, we have further modified the code to estimate the robustness of contextuality to dephasing noise. Our modification works for any strongly self-dual GPT with finite dimension, and the reasoning behind it is given in detail in Appendix B.
@@ I.3 Parity-oblivious multiplexing
Mathematically, one can define the parity of subsets of $x$ as follows: consider the parity bitstrings $\{t|t=(t_{1},...,t_{n}),\quad\sum_{i}t_{i}\geq 2\}$; then, $x\cdot t$ tell us the parity of some subset of $x$ determined by $t$ – and the condition $\sum t_{i}\geq 2$ ensures this subset has at least two bits. Therefore, the set $\{x\cdot t\}_{t}$ encodes the parity information of all subsets of $x$ containing more than one bit. With this notation, we say that a scenario satisfies parity-obliviousness if, given the parity strings $\{t\}$, it satisfies
@@
for all $t$, $x$ and $y$.
@@
In the quantum realisation of a $3$-to-1 parity-oblivious task that is going to be explored in this work, a 3-bitstring is encoded in different preparations of a qubit. Parity-obliviousness imposes that mixing preparations with equal parity must yield the same statistical mixture as mixing preparations with the opposite parity. This adds symmetries to the shape formed by the convex hull formed by the preparations. We exploit these symmetries by considering implementations whose set of states is given by a rectangular cuboid, with each vertex parameterized by the angle $\theta$ that all preparations equally form with respect to the $Z$ axis of the Bloch sphere. Measurements are assigned as the tomographic complete set $\hat{X},\hat{Y},\hat{Z}$ of Pauli measurements. This realisation obeys parity obliviousness constraints given in Eq. (4), and a schematic representation of this scenario can be found in Figure 2. Notice that even though Bob cannot retrieve more than $1$ bit of information in a single-shot, he can choose which information to retrieve (though not deterministically). This can only be done because, even though one can at most retrieve $1$ bit of information from a qubit, the qubit can carry more information than that. In this sense, Holevo’s theorem is obeyed, but communication advantage is also achieved.
@@
Figure 2: Representation of the preparation and measurement procedures of a 3-to-1 POM scenario among the family of scenarios investigated. Measurements lay on the $X$, $Y$, and $Z$ axes, while preparations are represented by the vertices of the hexahedron.
@@
In general, the resourcefulness of a POM task is usually quantified by the probability of getting the outcome for measurement $y$ that correctly matches the $y$-th bit in the string $x$, for any outcome and string, i.e.,
$s=\frac{1}{2^{n}n}\sum_{y=0}^{n}\sum_{x\in\{0,1\}^{\times n}}p(b=x_{y}|M_{y},P_{x}).$ (5)
@@
It has been demonstrated that indeed the success rate is a good quantifier for a resource theory of contextuality [46]. Moreover, it is known that quantum realisations of this protocol can exceed the noncontextual bound of $s_{NC}=\frac{1}{2}\left(1+\frac{1}{n}\right)$ [12], attesting that contextuality is the source of quantum advantage for these tasks, since quantum realisations can achieve success rates of up to $s_{Q}=\frac{1}{2}\left(1+\frac{1}{\sqrt{n}}\right)$ [14], constituting instances of quantum advantage for communication.

### v3

@@ Abstract
Generalised contextuality is the notion of nonclassicality powering up a myriad of quantum tasks, among which is the celebrated case of a two-party information processing task where classical information is compressed in a quantum channel, the parity-oblivious multiplexing (POM) task. The success rate is the standard quantifier of resourcefulness for this task, while robustness-based quantifiers are as operationally motivated and have known general properties. In this work, we leverage analytical and numerical tools to estimate robustness of contextuality in POM scenarios under different types of noise. We conclude that for the 3-to-1 case robustness of contextuality to depolarisation, as well as a minimisation of robustness of contextuality to dephasing over all bases, are good quantifiers for the nonclassical advantage of this scenario. Moreover, we obtain a general relation between robustness of contextuality to depolarisation and the success rate in any $n$-to-1 POM scenario and show how it can be used to bound the number of bits encoded in this task.
@@
Since the advent of quantum technologies, the superiority of quantum over classical methods has been anticipated across various areas, such as computation, information processing, and metrology. A celebrated result by Holevo [1] suggests that communication tasks may be an exception, stating that one can never retrieve more than $n$ bits of classical information out of $n$ qubits, apparently forbidding quantum superiority over classical strategies for communicating. Quantum advantage, however, can arise in communication tasks that profit from the quantum systems’ capacity to carry more information than classical ones [2, 3]. Therefore, such tasks prove interesting from both practical and foundational perspectives: they point to advantageous technological opportunities using quantum systems and also showcase how Holevo’s theorem can be conciliated with such an advantage in communication. We here focus on one of such tasks, called parity-oblivious multiplexing (POM), in which Alice sends Bob an $n$-bit string by sharing a smaller number of qubits. It has been demonstrated that the maximum success probability using quantum strategies for POM tasks is strictly higher than the maximum success probability obtained using classical systems, constituting an instance of quantum advantage over classical strategies for communication [4].
@@
It was previously shown that contextuality [5] powers quantum advantage in POM tasks [4]. Besides POM, contextuality powers up other aspects of communication [6, 7, 8], as well as computation [9], machine learning [10], information processing [11, 12, 13, 14], metrology [15], state-dependent cloning [16], and state discrimination [17, 18, 19, 20]. Contextuality first appeared in the literature in the Kochen-Specker theorem [21], and only recently it has been preliminary employed to make statements about generalised probabilistic theories (GPTs) [22, 23, 24] other than quantum theory. Generalisations of this notion have been developed in recent years, and Spekkens introduced an alternative approach that extends the notion of contextuality to encompass preparations, transformations, and unsharp measurements [5], contrasting with the Kochen-Specker theorem that primarily focuses on projective measurements. Since then, generalised contextuality has been proved to subsume or be related to many common notions of nonclassicality [4, 25, 26, 9, 27, 17, 28, 29, 30, 31, 32], while challenging the nonclassical status of some phenomena [33, 34, 35, 36, 37]. Within the generalised contextuality framework – and therefore throughout this manuscript – a theory or scenario therein is deemed classical when it can be explained by a generalised noncontextual ontological model.
@@
Since we have good and operationally motivated measures for contextuality (see Sec. II.2 for a discussion on this) – robustness against different kinds of noise – and we know that contextuality powers up the advantage behind parity-oblivious multiplexing tasks, it is natural to ask how these contextuality measures relate to the natural quantifier in POM tasks, i.e., the success probability itself. In this work, we turn to this problem by exploring the generality allowed by the code in Ref. [38] to define different kinds of noise and compare its behaviour to the success probability in POM tasks. We find a general expression relating the robustness of contextuality to depolarisation and the success rate for $n$-to-1 POM tasks and show how it can be used to recover a well-known bound on the number of bits optimally encoded in a qubit in this task. Then, building up on Ref. [40], we generalise its modification to estimate robustness of contextuality to dephasing noise in a family of scenarios related to the 3-to-1 POM. Our numerical results show that robustness of contextuality to dephasing can also perform as a quantifier when minimised over specific axes. Moreover, since the code essentially tests for simplex-embeddability of noisy states and effects, the method we adopt here can also be used for testing a class of generalised probabilistic theories broader than just quantum theory.
@@ I.1 Noncontextuality
The so-called operational approach is a way to construct physical theories based on laboratory experience [22, 23, 42]. This means that instead of the physical properties of a system and their time evolution, operational theories focus on how this physical system can be prepared, manipulated, and investigated in general experimental setups. Therefore, this approach is broad in the sense that it can be used to describe different kinds of theories, quantum and classical theories included. A typical prepare-and-measure scenario is described in this framework by the possible preparation procedures $P\in\mathcal{P}$, measurement procedures $M\in\mathcal{M}$, and possible outcomes $k\in K$ for each measurement that can be implemented in the experiment, as well as the statistics $p(k|M,P)$ obtained by implementing these procedures.
Moreover, these scenarios are also characterised by equivalences between different procedures: two preparations, denoted as $P$ and $P^{\prime}$, are deemed operationally equivalent, denoted as $P\simeq P^{\prime}$, if they cannot be distinguished empirically, even in principle, i.e.,
$P\simeq P^{\prime}\iff p(k|M,P)=p(k|M,P^{\prime}),\quad\forall k|M\in K\times\mathcal{M}.$ (1)
Similarly, two measurement outcomes, denoted as $k|M$ and $k^{\prime}|M^{\prime}$, are considered operationally equivalent, denoted as $k|M\simeq k^{\prime}|M^{\prime}$, if they both yield the same probability for any possible preparation, i.e.,
$k|M\simeq k^{\prime}|M^{\prime}\iff p(k|M,P)=p(k^{\prime}|M^{\prime},P),\quad\forall P\in\mathcal{P}.$ (2)
Because coarse-graining over measurement outcomes is an operationally valid measurement outcome, the notion of equivalent measurement procedures follows naturally from the equivalence between measurement outcomes. Notice that the equivalence relation $\simeq$ yields equivalence classes between the procedures, so that
$[P]:=\{P^{\prime}\in\mathcal{P}|P^{\prime}\simeq P\};$ (3)
The immediate interpretation is that the probability rule $p$ in the operational theory does not depend on the procedures $P$ and $k|M$ themselves, but only on the equivalence classes $[P]$ and $[k|M]$ they belong to. The excess information that singles out a particular procedure $P$ in the class $[P]$ (or a particular procedure $k|M$ in the class $[k|M]$) is what we refer to as a context. The tuple $(\mathcal{P},\mathcal{M},K,p,\simeq)$ defines an operational theory. As previously mentioned, quantum theory can be reformulated as one possible operational theory.
@@
Generalised contextuality (hereon referred to as contextuality since it is the relevant notion of contextuality for the purpose of this manuscript) arises within this operational framework as the impossibility of an operational theory to conform with a noncontextual ontological model (as we describe next). An ontological model for an operational theory provides an underlying explanation for its statistics based on classical probability theory and Boolean logic, such that the behaviours captured by the operational theory can be reasoned about in classical terms. In particular, the ontological model maps the system to an ontic space $\Lambda$, the elements $\lambda\in\Lambda$ are called ontic states, preparation procedures are associated with probability distributions $\mu_{P}(\lambda)$, called epistemic states, and measurement outcomes are associated with response functions222i.e., $\sum_{k\in K}\xi_{k|M}(\lambda)=1,\quad\forall M\in\mathcal{M},\lambda\in\Lambda.$ $\xi_{k|M}(\lambda)$ over the ontic space, with $\lambda\in\Lambda$. This model explains the operational theory when, for all $P\in\mathcal{P}$ and $k|M\in K\times\mathcal{M}$, it satisfies
@@
Admitting of some ontological model is not yet a classical feature but a general one. Indeed, classicality has additional demands: any equivalences between operational procedures must hold for their representation using ontological models [^3] Otherwise, further arguments must be provided on why the procedures are not operationally distinguishable, despite being ontologically different., which underpins the notion of noncontextuality [5, 43]. An ontological model satisfies the assumption of noncontextuality if, for any $P,P^{\prime}\in\mathcal{P}$, $k|M,k^{\prime}|M^{\prime}\in K\times\mathcal{M}$,
@@
In other words, an ontological model is noncontextual if the contexts are not relevant for the ontological explanation underlying the experiment whenever the statistics of the operational theory do not depend on those contexts. An ontological model satisfying this constraint is said noncontextual, and a theory is contextual when it does not admit such a description.
The condition of admitting of some noncontextual ontological model has proven to be a precise definition of classicality (when classicality is understood as the theory admiting of an embedding into a simplicial theory [39]) since the noncontextual ontological model itself provides a classical explanation to the empirical predictions. Hence whenever we say classical throughout this article, we mean generalised-noncontextual. As was shown before, quantum theory does not admit such explanations [5], featuring nonclassicality.
@@ I.2 Operational measures of contextuality
Often, assessing noncontextuality of an operational scenario requires the characterisation of a large ontic space [^4] In quantum scenarios, for instance, the ontic space might be as large as $|\Lambda|=(\text{dim}\mathcal{H})^{2}$ [38]. and the possible epistemic states and response functions over it that can explain the statistical data and conform to the assumption of noncontextuality for all equivalence relations. To avoid these issues when assessing the nonclassicality of the examples investigated in this work, we will employ the linear program introduced in Ref. [38], which relies on results from the framework of generalised probabilistic theories (GPT) [22, 23, 24]. We will not enter into the details of how the operational scenario translates to the GPT description, but a summary of the linear program and the reasoning behind it is provided in A. The relevant aspect for the present work is that this linear program decides the existence of a simplex embedding for the GPT fragment (see Fig. 1), which in turn is equivalent to the existence of a noncontextual ontological model for the associated operational scenario [39]. If the program cannot find a simplex embedding for the input scenario, i.e., the scenario exhibits contextuality, it calculates how much depolarising noise $r$ is necessary so that the simplex embedding becomes possible. The implementation of the program takes in sets of states and effects of a GPT and outputs this minimal value of $r$ along with the respective noncontextual ontological model explaining the scenario partially depolarised by a factor $r$.
@@
The quantity $r$ hence may serve as an operational measure of nonclassicality, which we denote as the robustness of contextuality to depolarisation. It is known from simplex embedding reasoning that there is always a finite amount of partial depolarising noise under which an operational scenario becomes noncontextual, and explicit examples of this fact have been reported [17, 44, 45]. Alternatively, robustness can be defined in terms of different kinds of noise besides depolarising, so in this work, we also consider the case of dephasing. This is because it is a well-established fact that a completely dephased prepare-and-measure scenario will always admit of a noncontextual model [40], so the action of dephasing noise over a scenario will also eventually allow for a simplex embedding [^5] Although in this case, partial dephasing noise might not be enough to allow for the simplex embedding, as shown in Ref. [40].. Therefore, we have further modified the code to estimate the robustness of contextuality to dephasing noise. Our modification works for any strongly self-dual GPT with finite dimension, and the reasoning behind it is given in detail in B.
@@ I.3 Parity-oblivious multiplexing
Mathematically, one can define the parity of subsets of $x$ as follows: consider the parity bitstrings $\{t|t=(t_{1},...,t_{n}),\quad\sum_{i}t_{i}\geq 2\}$; then, $x\cdot t$ tell us the parity of some subset of $x$ determined by $t$ – and the condition $\sum t_{i}\geq 2$ ensures this subset has at least two bits. Therefore, the set $\{x\cdot t:=\sum_{i}x_{i}t_{i}\}_{t}$ encodes the parity information of all subsets of $x$ containing more than one bit. With this notation, we say that a scenario satisfies parity-obliviousness if, given the parity strings $\{t\}$, it satisfies
@@
for all $t$, $x$ and $y$. This is why generalised contextuality is of key interest for this protocol: for the $n$-to-1 POM there are $2^{n}$ possible preparations that must all be indistinguishable once parity about the encoded bitstring is lost. Kochen-Specker contextuality [21], for instance, cannot give an account of scenarios with multiple preparations and could only assess the nonclassicality related to the task of retriving bits of a specified encoded bitstring, something that displays no quantum advantage according to Holevo’s result.
@@
In the quantum realisation of a $3$-to-1 parity-oblivious task that is going to be explored in this work, a 3-bitstring is encoded in different preparations of a qubit. Parity-obliviousness imposes that mixing preparations with equal parity must yield the same statistical mixture as mixing preparations with the opposite parity. This adds symmetries to the shape formed by the convex hull formed by the preparations. We exploit these symmetries by considering implementations whose set of states is given by a rectangular cuboid, with each vertex parameterized by the angle $\theta$ that all preparations equally form with respect to the $Z$ axis of the Bloch sphere. Measurements are assigned as the tomographic complete set $\hat{X},\hat{Y},\hat{Z}$ of Pauli measurements. This realisation obeys parity obliviousness constraints given in Eq. (8), and a schematic representation of this scenario can be found in Figure 2, where the possible states are given explicitly by
$\rho_{(x_{1},x_{2},x_{3})}$ $=$ $\frac{1}{2}\left(\mathbb{1}+\frac{(-1)^{x_{1}}}{\sqrt{2}}\sin\theta\hat{X}+\frac{(-1)^{x_{2}}}{\sqrt{2}}\sin\theta\hat{Y}\right.$ (9)
$\qquad\qquad\qquad\qquad\left.+(-1)^{x_{3}}\cos\theta\hat{Z}\right),$
for $\theta\in\left[0,\frac{\pi}{2}\right]$, while the measurement outcomes are projections onto the eigenvectors of the Pauli matrices, i.e.,
$E_{k|M_{y}}:=\frac{1}{2}(\mathbb{1}+(-1)^{k}\hat{M}_{y}),$ (10)
where $k\in\{0,1\}$ and $\hat{M}_{1}:=\hat{X}$, $\hat{M}_{2}:=\hat{Y}$, and $\hat{M}_{3}:=\hat{Z}$ are the Pauli operators.
@@
Figure 2: Representation of the preparation procedures of a 3-to-1 POM scenario among the family of scenarios investigated. Measurements lie on the $X$, $Y$, and $Z$ axes, while preparations are represented by the vertices of the hexahedron, lying on thee surface of the Bloch sphere. Each of the 8 possible bitstrings is encoded in a different vertex of the polytope, and fixing a value for the angle $\theta$ between all states and the $Z$ axis singles out a POM scenario.
Notice that even though Bob cannot retrieve more than $1$ bit of information in a single-shot, he can choose which information to retrieve (though not deterministically). This can only be done because, even though one can at most retrieve $1$ bit of information from a qubit, the qubit can carry more information than that. In this sense, Holevo’s theorem is obeyed, but communication advantage is also achieved.
@@
In general, the resourcefulness of a POM task is quantified by the probability of getting the outcome for measurement $y$ that correctly matches the $y$-th bit in the string $x$, for any outcome and string, i.e.,
$s=\frac{1}{2^{n}n}\sum^{n}_{y=1}\,\sum_{x\in X_{n}}p(k=x_{y}|M_{y},P_{x}),$ (11)
@@
where $X_{n}:=\{0,1\}^{\times n}$ is the set of all possible bitstrings that can be encoded in the task. It has been demonstrated that indeed the success rate is a good quantifier for a resource theory of contextuality [46]. Moreover, it is known that quantum realisations of this protocol can exceed the noncontextual bound of $s_{NC}=\frac{1}{2}\left(1+\frac{1}{n}\right)$ [11], attesting that contextuality is the source of quantum advantage for these tasks, since quantum realisations can achieve success rates of up to $s_{Q}=\frac{1}{2}\left(1+\frac{1}{\sqrt{n}}\right)$ [13], constituting instances of quantum advantage for communication.